# EventFlow: Forecasting Continuous-Time Event Data with Flow Matching

### v1

@@ Abstract
Continuous-time event sequences, in which events occur at irregular intervals, are ubiquitous across a wide range of industrial and scientific domains. The contemporary modeling paradigm is to treat such data as realizations of a temporal point process, and in machine learning it is common to model temporal point processes in an autoregressive fashion using a neural network. While autoregressive models are successful in predicting the time of a single subsequent event, their performance can be unsatisfactory in forecasting longer horizons due to cascading errors. We propose EventFlow, a non-autoregressive generative model for temporal point processes. Our model builds on the flow matching framework in order to directly learn joint distributions over event times, side-stepping the autoregressive process. EventFlow is likelihood-free, easy to implement and sample from, and either matches or surpasses the performance of state-of-the-art models in both unconditional and conditional generation tasks on a set of standard benchmarks.
@@ 1 Introduction
Many stochastic processes, ranging from consumer behavior (Hernandez et al., 2017) to the occurrence of earthquakes (Ogata, 1998), are best understood as a sequence of discrete events which occur at random times. Any observed event sequence, consisting of one or more event times, may be viewed as a draw from a temporal point process (TPP) (Daley & Vere-Jones, 2003) which characterizes the distribution over such sequences. Given a collection of observed event sequences, faithfully modeling the underlying TPP is critical in both understanding and forecasting the phenomenon of interest.
While multiple different parametric TPP models have been proposed (Hawkes, 1971; Isham & Westcott, 1979), their limited flexibility limits their application when modeling complex real-world sequences. This has motivated the use of neural networks (Du et al., 2016; Mei & Eisner, 2017) in modeling TPPs. To date, most neural network based TPP models are autoregressive in nature (Shchur et al., 2020a; Zhang et al., 2020), where a model is trained to predict the next event time given an observed history of events. However, in many tasks, we are interested not only in the next event, but in the entire sequence of events which is to follow. While these models can achieve high likelihoods, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
@@
Motivated by these limitations, we propose EventFlow, a generative model which directly learns the joint event time distributions, thus allowing us to avoid autoregressive sampling altogether. Our proposed model extends the flow matching framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to the setting of TPPs, where we learn a continuous flow from a reference TPP to our data TPP. At an intuitive level, samples from our model are generated by drawing a collection of event times from a reference distribution and flowing these events along a learned vector field. The number of events is fixed throughout this process, decoupling the event counts and their times, so that the distribution over event counts can be learned or otherwise specified. See Figure 1 for an illustration. As our primary contribution regards the modeling of the event times themselves, we focus on unmarked point processes in this work. More specifically,
@@
  • Our model provides a new perspective on modeling TPPs and sidesteps common pitfalls in existing approaches. In particular, EventFlow is likelihood-free, non-autoregressive, easy to sample from, and straightforward to implement.
  • On standard benchmark datasets, EventFlow obtains strong performance on a multi-step forecasting task, outperforming the second-best model by a factor of 2.2 to 4, and matches or exceeds the performance of state-of-the-art models for unconditional generation.
@@ 2 Related Work
Numerous models have been proposed for both components. Popular choices for the history encoder include RNN-based models (Du et al., 2016; Shchur et al., 2020a; Mei et al., 2019) or attention-based models (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022). While attention-based encoders can provide longer-range contexts, this benefit typically comes at the cost of additional memory overhead. Similarly, a wide range of forms for the event decoder have also been proposed. The most common approach is to parametrize a conditional intensity function via a neural network.
For instance, several authors (Mei & Eisner, 2017; Zuo et al., 2020; Zhang et al., 2020) model the conditional intensity using a parametric form inspired by the Hawkes process (Hawkes, 1971), and Du et al. (2016) model the (log-)conditional intensity through an affine function of the history embedding. Similarly, Okawa et al. (2019) model the conditional intensity using a mixture of Gaussian kernels.
Most closely related to our work are approaches which use generative models as decoders. These models often do not assume a parametric form for the decoder, enhancing their flexibility. For instance, Xiao et al. (2017b) propose the use of W-GANs to generate new events. Similarly, Shchur et al. (2020a) learn the distribution over the next inter-arrival time via a normalizing flow. Lin et al. (2022) benchmark several choices of generative models, including diffusion, GANs, and VAEs. Despite the flexibility of these models, these approaches are all autoregressive in nature, making them ill-suited for multi-step forecasting tasks. In contrast, Lüdke et al. (2023) propose a diffusion-style model which is able to avoid autoregressive sampling via an iterative refinement procedure.
Our work can be viewed as a novel approach for building flexible decoders for TPPs, extending flow matching to the setting of continuous-time event sequences. In contrast to prior work using generative models, our model is likelihood-free and non-autoregressive, achieving strong performance on long-term forecasting tasks. The work of Lüdke et al. (2023) is perhaps most closely related to ours, but we emphasize that the method of Lüdke et al. (2023) requires an involved training and sampling procedure. In contrast, our method is straightforward to both implement and sample from, while simultaneously outperforming existing approaches.
@@ 3 Autoregressive TPP Models
In the first approach, a conditional probability density of the form $p(t\mid\mathcal{H}_{t_{n}})$ is learned, allowing us to specify a joint distribution over event times $p(t_{1},\dots,t_{n})$ autoregressively via ${p(t_{1},\dots,t_{n})=p(t_{1})\prod_{k=2}^{n}p(t_{k}\mid\mathcal{H}_{t_{k-1}})}$.
Alternatively, we may define the conditional intensity $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t^{-}})=p(t\mid\mathcal{H}_{t_{n}})/(1-F(t\mid\mathcal{H}_{t_{n}}))$, where $F(t\mid\mathcal{H}_{t_{n}})=\int_{t_{n}}^{t}p(s\mid\mathcal{H}_{t_{n}})\,\mathrm{d}s$ is the CDF associated with $p(t\mid\mathcal{H}_{t_{n}})$. Informally, the conditional intensity $\lambda^{*}(t)$ can be thought of (Rasmussen, 2011) as the instantaneous rate of occurrence of events at time $t$ given the previous $n$ events and given that no events have occurred since $t_{n}$. By integrating $\lambda^{*}(t)$, one can show that
@@
$L\left(\{t_{k}\}\right)=p(t_{1},\dots,t_{n})\left(1-F(T\mid\mathcal{H}_{t_{n}})\right)=\left(\prod_{k=1}^{n}\lambda^{*}(t_{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (2)
where the CDF term is included to indicate that no events beyond $t_{n}$ have occurred and $\Lambda^{*}(T)=\int_{0}^{T}\lambda^{*}(s)\,\mathrm{d}s$ is the total intensity. Autoregressive models are typically trained by maximizing this likelihood (Du et al., 2016; Mei & Eisner, 2017; Shchur et al., 2020a). We emphasize that this likelihood is not simply the joint event-time density $p_{n}(t_{1},\dots,t_{n})$, as the likelihood measures the fact that no events occur after $t_{n}$.
It is worth noting that evaluating $L(\{t_{k}\})$ can be non-trivial in practice. For models which parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei & Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring several forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$. Models which directly parametrize the density $p(t\mid\mathcal{H}_{t})$ suffer from the same drawback when computing the corresponding CDF in Equation (2). Moreover, some approaches, such as the diffusion-based approach of Lin et al. (2022), are only trained to maximize an ELBO of $p(t\mid t_{1},\dots,t_{n})$, and are thus unable to compute the proper likelihood in Equation (2).
@@
In many tasks, we are interested not only in an accurate model of the intensity (or distribution), but also sampling new event sequences from the corresponding distribution. For instance, when forecasting an event sequence, we may want to generate several forecasts in order to provide uncertainty quantification over these predictions. Stochastic forecasts also have the potential to enable probabilistic queries (Boyd et al., 2023; 2022), e.g., estimating hitting times (the time it will take until a specified number of events occurs) or estimating the number of events in some subwindow. However, sampling from existing autoregressive models can be difficult.
For instance, the flow-based model of Shchur et al. (2020a) requires a numerical approximation to the inverse of the model to perform sampling. Similarly, the diffusion-based approach of Lin et al. (2022) can require several hundred forward passes of the model to generate a single event time, rendering it costly when generating long sequences. Moreover, the predictive performance of autoregressive models is often unsatisfactory on multi-step generation tasks due to the accumulation of errors over many steps (Lin et al., 2021; Lüdke et al., 2023). This difficulty is particularly pronounced for intensity-based models (Du et al., 2016; Mei & Eisner, 2017; Zhang et al., 2020), where naively computing the implied distribution in Equation (1) is prohibitively expensive. Instead, sampling from intensity-based models is typically achieved via the thinning algorithm (Ogata, 1981; Lewis & Shedler, 1979). However, this algorithm has several hyperparameters to tune, is challenging to parallelize, and can be difficult for practitioners to implement (Xue et al., 2024).
@@ 4 EventFlow
Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages over existing approaches. First, EventFlow directly models the joint distribution over event times, thereby avoiding autoregression entirely. Second, our model is likelihood-free, avoiding the Monte Carlo estimates needed to estimate the likelihood in Equation (2) during training. Third, sampling from our model amounts to solving an ordinary differential equation. This is straightforward to implement and parallelize, allowing us to avoid the difficulties of thinning-based approaches used in existing models. We build upon the flow matching (or stochastic interpolant) framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to develop our model. We begin below by focusing on the unconditional setting, and later discuss how to extend our method for conditional generation as necessary.
4.1 Preliminaries
@@
We first introduce some necessary background and notation. Let $\mathcal{T}=[0,T]$ be a finite-length interval. The set $\Gamma$ denotes the configuration space of $\mathcal{T}$ (Albeverio et al., 1998), i.e., the set of all finite counting measures on the set $[0,T]$. A point $\gamma\in\Gamma$ corresponds to a measure of the form $\gamma=\sum_{k=1}^{n}\delta[t^{k}]$, i.e., a finite collection of Dirac deltas located at event times $t^{k}\in\mathcal{T}$. A temporal point process (TPP) on $\mathcal{T}$ is a probability distribution $\mu$ over the configuration space $\Gamma$. Informally, $\mu$ represents a distribution over sequences $\gamma$ living in the configuration space $\Gamma$ which constitutes the set of valid sequences. We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the TPP realization $\gamma$ [^1] This can be thoughts of in terms of the counting process, i.e. $N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$. While it is common to represent TPPs as distributions over random sets of event times, in our approach it will be more convenient to represent TPPs as random measures (Kallenberg et al., 2017).
@@
We assume all TPP distributions are atomless (Kallenberg et al., 2017, Ch. 1), i.e., the probability of observing an event at any singleton is zero. In addition, we assume all TPPs are simple (Kallenberg et al., 2017, Ch. 2), i.e., no more than one event can occur simultaneously. Under these assumptions, a TPP $\mu$ can be fully characterized (Daley & Vere-Jones, 2003, Prop. 5.3.II) by a probability distribution which specifies the number of events and a collection of joint densities corresponding to the event times themselves. In a slight abuse of notation, we will write $\mu(n)$ for the corresponding distribution over event counts, and $\{\mu^{n}(t^{1},\dots,t^{n})\}_{n=1}^{\infty}$ for the collection of joint distributions. In other words, for any given $n\in\mathbb{Z}_{\geq 0}$, the probability of observing $n$ events in the interval $\mathcal{T}$ is $\mu(n)$, and $\mu^{n}(t^{1},\dots,t^{n})$ describes the corresponding joint distribution of event times. We further restrict each $\mu^{n}$ to be supported only on the ordered sets, so that we may assume $t^{1}<t^{2}<\dots<t^{n}$.
@@
4.2 Balanced Couplings
Our first step is to define a useful notion of couplings (Villani et al., 2009), allowing us to pair event sequences drawn from $\mu_{0}$ with those drawn from $\mu_{1}$. A coupling between two TPPs $\mu,\nu\in\mathbb{P}(\Gamma)$ is a joint probability measure $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ over pairs of event sequences $(\gamma_{0},\gamma_{1})$ such that the marginal distributions of $\rho$ are $\mu$ and $\nu$. We say that the coupling $\rho$ is balanced if draws $(\gamma_{0},\gamma_{1})\sim\rho$ are such that $N(\gamma_{0})=N(\gamma_{1})$ almost surely. In other words, balanced couplings only pair event sequences with equal numbers of events. We will use $\Pi_{b}(\mu,\nu)$ to denote the set of balanced couplings of $\mu,\nu$. We can show that the set $\Pi_{b}(\mu,\nu)$ is nonempty if and only if the event count distributions of $\mu$ and $\nu$ are equal, placing a structural constraint on the suitable choices of a reference measure.
@@ 4.2 Balanced Couplings
We provide a proof in Appendix B. To provide some intuition for the proof, in one direction, we may explicitly construct a balanced coupling (which we describe in the next paragraph). Conversely, if a balanced coupling exists, the event count distribution of $\mu$ (and hence $\nu$) may be obtained by sampling from the coupling and counting the number of events occurring in either coordinate.
In practice, we follow a simple strategy for choosing both the reference TPP $\mu_{0}$ and the coupling $\rho$. Suppose $q$ is any given density on our state space $\mathcal{T}$, e.g., a uniform distribution. We take $\mu_{0}$ to be a mixed binomial process (Kallenberg et al., 2017, Ch. 3) whose event count distribution is given by that of the data $\mu_{1}(n)$ and joint event distributions given by independent products of $q$ (up to sorting). That is, to sample from $\mu_{0}$, one can simply sample $n\sim\mu_{1}(n)$ from the empirical event count distribution followed by sampling and sorting $n$ i.i.d. points $t^{k}\sim q$.
To draw a sample from our coupling $\rho$, we first sample a data sequence $\gamma_{1}\sim\mu_{1}$, followed by sampling $N(\gamma_{1})$ events independently from $q$ and sorting to produce a draw $\gamma_{0}\sim\mu_{0}$. We call this coupling the independent balanced coupling of $\mu$ and $\nu$.
4.3 Interpolant Construction
We now proceed to construct our interpolant $\eta_{s}\in\mathbb{P}(\Gamma)$. We will construct this path of TPPs via a local procedure which we then marginalize over a given balanced coupling. To that end, let $\rho$ be any balanced coupling of the reference measure $\mu_{0}$ and the data measure $\mu_{1}$, and suppose $z:=(\gamma_{0},\gamma_{1})\sim\rho$ is a draw from this coupling. As $\rho$ is balanced, we have that
$\gamma_{0}=\sum_{k=1}^{n}\delta[t_{0}^{k}]\qquad\gamma_{1}=\sum_{k=1}^{n}\delta[t_{1}^{k}]$ (3)
are both a collection of $n$ events. As TPPs are fully characterized by their joint distributions over event times, we will henceforth describe our procedure for a fixed (but arbitrary) number of events $n$. First, we the define measure $\gamma_{s}^{z}\in\Gamma$ via
@@ 4.3 Interpolant Construction
In order to perform the marginalization step, we now lift this deterministic path $(\gamma_{s}^{z})_{s=0}^{1}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s=0}^{1}\subset\mathbb{P}(\Gamma)$. We define the point process distribution $\eta_{s}^{z}\in\mathbb{P}(\Gamma)$ implicitly by adding independent Gaussian noise to each of the events in $\gamma_{s}^{z}$. That is, a draw $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ may be simulated via
@@
Finally, for any $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}$ via $\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$. Observe that, by construction, the event count distribution $\eta_{s}(n)$ is given by $\mu_{1}(n)$ for all $s\in[0,1]$. This path of TPP distributions $\eta_{s}$ approximately interpolates from the reference TPP $\mu_{0}$ at $s=0$ to the data TPP $\mu_{1}$ at $s=1$, in the sense that at the endpoints, the joint event time distributions $\eta_{0}^{n}(t^{1},\dots,t^{n})$ and $\eta_{1}^{n}(t^{1},\dots,t^{n})$ are given by a convolution of $\mu_{0}^{n}(t^{1},\dots,t^{n})$ and $\mu_{1}^{n}(t^{1},\dots,t^{n})$ with the Gaussian $\mathcal{N}(0,\sigma^{2}I_{n})$. As $\sigma^{2}\downarrow 0$, it is clear that we recover a genuine interpolant (Tong et al., 2024).
We now shift our attention to the transport of a single event $t_{s}^{k}$ for a fixed $k$. Through the addition of Gaussian noise, we have constructed a path of Gaussian distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$ whose mean is determined by the location of the $k$th event at the flow time $s$. This transport of a Gaussian can be achieved infinitesimally through the constant vector field ${v_{s}^{k}:[0,T]\to\mathbb{R}}$ given by $v_{s}^{k}(t)=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024). Thus, the evolution in (5) is generated by the vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ given by

### v2

# EventFlow: Forecasting Temporal Point Processes with Flow Matching

@@ Abstract
Continuous-time event sequences, in which events occur at irregular intervals, are ubiquitous across a wide range of industrial and scientific domains. The contemporary modeling paradigm is to treat such data as realizations of a temporal point process, and in machine learning it is common to model temporal point processes in an autoregressive fashion using a neural network. While autoregressive models are successful in predicting the time of a single subsequent event, their performance can degrade when forecasting longer horizons due to cascading errors and myopic predictions. We propose EventFlow, a non-autoregressive generative model for temporal point processes. The model builds on the flow matching framework in order to directly learn joint distributions over event times, side-stepping the autoregressive process. EventFlow is simple to implement and achieves a 20%-53% lower error than the nearest baseline on standard TPP benchmarks while simultaneously using fewer model calls at sampling time.
@@ 1 Introduction
Many stochastic processes, ranging from the occurrence of earthquakes (Ogata, 1998) to consumer behavior (Xu et al., 2014), are best understood as a sequence of discrete events which occur at random times. Any observed event sequence, consisting of one or more event times, may be viewed as a draw from a temporal point process (TPP) (Daley & Vere-Jones, 2003) which characterizes the distribution over such sequences. Given a collection of observed event sequences, faithfully modeling the underlying TPP is critical in both understanding and forecasting the phenomenon of interest.
While multiple different parametric TPP models have been proposed (Hawkes, 1971; Isham & Westcott, 1979), their limited flexibility limits their application when modeling complex real-world sequences. This has motivated the use of neural networks (Du et al., 2016; Mei & Eisner, 2017) in modeling TPPs. To date, most neural network based TPP models are autoregressive in nature (Shchur et al., 2020a; Zhang et al., 2020), where a model is trained to predict the next event time given an observed history of events. However, in many tasks, we are interested not only in the next event, but in the entire sequence of events which is to follow. Despite their popularity, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
@@
Motivated by these limitations, we propose EventFlow, a generative model which directly learns the joint event time distributions, thus allowing us to avoid autoregressive sampling altogether. Our proposed model extends the flow matching framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to the setting of TPPs, where we learn a continuous flow from a reference TPP to our data TPP. At an intuitive level, samples from our model are generated by drawing a collection of event times from a reference distribution and flowing these events along a learned vector field. The number of events is fixed throughout this process, decoupling the event counts and their times, so that the distribution over event counts can be learned or otherwise specified. See Figure 1 for an illustration. More specifically,
@@
  • We provide new perspective on modeling TPPs which sidesteps common pitfalls. In particular, the key idea of EventFlow is to decompose the generative process into a learned event count distribution and a flow-based generative model for the joint distribution of event times. Our model is non-autoregressive, easy to sample from, and straightforward to implement.
  • On standard benchmarks, we demonstrate that EventFlow obtains state-of-the-art performance on a multi-step forecasting task and matches or exceeds the performance of state-of-the-art models for unconditional generation. In particular, we achieve a 20%-53% reduction in forecasting error from the nearest baseline. Compared with existing methods, which can require hundreds of model evaluations at sampling time, we demonstrate that EventFlow can achieve competitive performance with only a single forward pass.
@@ 2 Related Work
Numerous models have been proposed for both components. Popular choices for the history encoder include RNN-based models (Du et al., 2016; Shchur et al., 2020a; Mei et al., 2019) or attention-based models (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022). While attention-based encoders can provide longer-range contexts, this benefit typically comes at the cost of additional memory overhead.
Similarly, a wide range of forms for the event decoder have also been proposed. The most common approach is to parametrize a conditional intensity function via a neural network (Mei & Eisner, 2017; Zuo et al., 2020; Du et al., 2016). An alternative approach is to use generative models as decoders. These models often do not assume a parametric form for the decoder, enhancing their flexibility. For instance, Xiao et al. (2017b) propose the use of W-GANs to generate new events. Similarly, Shchur et al. (2020a) learn the distribution over the next inter-arrival time via a normalizing flow. Lin et al. (2022) benchmark several choices of generative models, including diffusion, GANs, and VAEs. Despite the flexibility of these models, these approaches are all autoregressive in nature, making them ill-suited for multi-step forecasting tasks. In contrast, Lüdke et al. (2023) propose a diffusion-style model which is able to avoid autoregressive sampling via an iterative refinement procedure.
Our work can be viewed as a novel approach to building flexible decoders for TPPs, extending flow matching to the setting of continuous-time event sequences. In contrast previous generative approaches, our model is trained in a likelihood-free manner and is non-autoregressive. The work of Lüdke et al. (2023) is most closely related to ours, but we emphasize that Lüdke et al. (2023) require an involved training and sampling procedure. In contrast, our method is straightforward to both implement and sample from, while simultaneously outperforming existing approaches.
@@ 3 Autoregressive TPP Models
In the first approach, a conditional probability density of the form $p_{n}(t\mid\mathcal{H}_{t_{n}})$ is learned, allowing us to specify a joint distribution over event times $p(t_{1},\dots,t_{n})$ autoregressively via ${p_{n}(t_{1},\dots,t_{n})=p(t_{1})\prod_{k=2}^{n}p(t_{k}\mid\mathcal{H}_{t_{k-1}})}$. Alternatively, we may define the conditional intensity $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t^{-}})=p(t\mid\mathcal{H}_{t_{n}})/(1-F(t\mid\mathcal{H}_{t_{n}}))$, where $F(t\mid\mathcal{H}_{t_{n}})=\int_{t_{n}}^{t}p(s\mid\mathcal{H}_{t_{n}})\,\mathrm{d}s$ is the CDF associated with $p(t\mid\mathcal{H}_{t_{n}})$. Informally, the conditional intensity $\lambda^{*}(t)$ can be thought of (Rasmussen, 2011) as the instantaneous rate of occurrence of events at time $t$ given the previous $n$ events and given that no events have occurred since $t_{n}$. By integrating $\lambda^{*}(t)$, one can show that
@@
$L\left(\{t_{k}\}_{k=1}^{n}\right)=p_{n}(t_{1},\dots,t_{n})\left(1-F(T\mid\mathcal{H}_{t_{n}})\right)=\left(\prod_{k=1}^{n}\lambda^{*}(t_{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (2)
where the CDF term is included to indicate that no events beyond $t_{n}$ have occurred and $\Lambda^{*}(T)=\int_{0}^{T}\lambda^{*}(s)\,\mathrm{d}s$ is the total intensity. Autoregressive models are typically trained by maximizing this likelihood (Du et al., 2016; Mei & Eisner, 2017; Shchur et al., 2020a). We emphasize that this likelihood is not simply the joint event-time density $p_{n}(t_{1},\dots,t_{n})$, as the likelihood also measures the fact that no events occur between $t_{n}$ and $T$.
It is worth noting that evaluating $L(\{t_{k}\})$ can be non-trivial in practice. For models that parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei & Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring many forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$. Models which directly parametrize the density $p(t\mid\mathcal{H}_{t})$ suffer from the same drawback when computing the corresponding CDF in Equation (2). Moreover, some approaches, such as the diffusion-based approach of Lin et al. (2022), are only trained to maximize an ELBO of $p(t_{k}\mid t_{1},\dots,t_{k-1})$, and are thus unable to compute the proper likelihood in Equation (2).
@@
In many tasks, we are interested not only in an accurate model of the intensity (or distribution), but also sampling new event sequences. For instance, when forecasting an event sequence, we may want to generate several forecasts in order to provide uncertainty quantification over these predictions. However, sampling from existing models can be difficult. The flow-based model of Shchur et al. (2020a) requires a numerical approximation to the inverse of the model to perform sampling. The diffusion-based approach of Lin et al. (2022) can require several hundred forward passes of the model to generate a single event time, rendering it costly when generating long sequences. Moreover, the predictive performance of autoregressive models is often unsatisfactory on multi-step generation tasks due to the accumulation of errors over many steps (Lin et al., 2021; Lüdke et al., 2023). This difficulty is particularly pronounced for intensity-based models (Du et al., 2016; Mei & Eisner, 2017; Zhang et al., 2020), where naively computing the implied distribution in Equation (1) is prohibitively expensive. Instead, sampling from intensity-based models is typically achieved via the thinning algorithm (Ogata, 1981; Lewis & Shedler, 1979). However, this algorithm has several hyperparameters to tune, is challenging to parallelize, and can be difficult for practitioners to implement (Xue et al., 2024). For instance, the thinning algorithm requires one to know an upper bound on the intensity, which is in practice selected heuristically.
@@ 4 EventFlow
Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages. First, EventFlow directly models the joint distribution over event times, thereby avoiding autoregression entirely. Second, our model is likelihood-free at training time, avoiding the Monte Carlo estimates needed to estimate the likelihood in Equation (2). Third, sampling from our model amounts to solving an ordinary differential equation. This is straightforward to implement and parallelize, allowing us to avoid the difficulties of thinning-based approaches. In fact, EventFlow can achieve competitive performance using only a single forward pass at sampling time. We build upon the flow matching (or stochastic interpolant) framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to develop our model. We begin below by focusing on the unconditional setting, followed by an extension for conditional generation.
Preliminaries
@@
We first introduce some necessary background and notation. Let $\mathcal{T}=[0,T]$ be a finite length interval. The set $\Gamma$ denotes the configuration space of $\mathcal{T}$ (Albeverio et al., 1998), i.e., the set of all finite counting measures on the set $[0,T]$. A point $\gamma\in\Gamma$ corresponds to a measure of the form $\gamma=\sum_{k=1}^{n}\delta[t^{k}]$, i.e., a finite collection of Dirac deltas located at event times $t^{k}\in\mathcal{T}$. A temporal point process (TPP) on $\mathcal{T}$ is a probability distribution $\mu$ over the configuration space $\Gamma$. Informally, $\mu$ represents a distribution over finite sequences $\gamma\in\Gamma$ whose events are in $\mathcal{T}$. We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the sequence $\gamma$ [^1] This can be thought of in terms of the counting process, i.e., $N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$. While it is common to represent TPPs as distributions over random sets of event times, in our approach it will be more convenient to represent TPPs as random measures (Kallenberg, 2017).
@@
Under mild assumptions, a TPP $\mu$ can be fully characterized (Daley & Vere-Jones, 2003, Prop. 5.3.II) by a probability distribution which specifies the number of events and a collection of joint densities corresponding to the event times themselves. In a slight abuse of notation, we will write $\mu(n)$ for the corresponding distribution over event counts, and $\{\mu^{n}(t^{1},\dots,t^{n})\}_{n=1}^{\infty}$ for the collection of joint distributions. In other words, for any given $n\in\mathbb{Z}_{\geq 0}$, the probability of observing $n$ events in the interval $\mathcal{T}$ is $\mu(n)$, and $\mu^{n}(t^{1},\dots,t^{n})$ describes the corresponding joint distribution of event times. We further restrict each $\mu^{n}$ to be supported only on the ordered sets, so that $t^{1}<t^{2}<\dots<t^{n}$.
@@
Balanced Couplings
Our first step is to define a useful notion of couplings, allowing us to pair event sequences drawn from $\mu_{0}$ with those drawn from $\mu_{1}$. A coupling between two TPPs $\mu,\nu\in\mathbb{P}(\Gamma)$ is a joint probability measure $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ over pairs of event sequences $(\gamma_{0},\gamma_{1})$ such that the marginal distributions of $\rho$ are $\mu$ and $\nu$. We say that the coupling $\rho$ is balanced if draws $(\gamma_{0},\gamma_{1})\sim\rho$ are such that $N(\gamma_{0})=N(\gamma_{1})$ almost surely. In other words, balanced couplings only pair event sequences with equal numbers of events. While we will later see how to interpolate between any two given event sequences, this coupling will allow us to decide which sequences to interpolate. In particular, a balanced coupling will allow us to pair sequences such that they always have the same number of events, allowing us to avoid the addition or deletion of events during both training and sampling. We will use $\Pi_{b}(\mu,\nu)$ to denote the set of balanced couplings. The following proposition shows $\Pi_{b}(\mu,\nu)$ is nonempty if and only if the event count distributions of $\mu$ and $\nu$ are equal, placing a structural constraint on the suitable choices of $\mu_{0}$.
@@
We provide a proof in Appendix B. In practice, we follow a simple strategy for choosing both the reference TPP $\mu_{0}$ and the coupling $\rho$. Suppose $q$ is any given density on our state space $\mathcal{T}$, e.g., a uniform distribution. We take $\mu_{0}$ to be a mixed binomial process (Kallenberg, 2017, Ch. 3) whose event count distribution is given by that of the data $\mu_{1}(n)$ and joint event distributions given by independent products of $q$ (up to sorting). That is, to sample from $\mu_{0}$, one can simply sample $n\sim\mu_{1}(n)$ from the empirical event count distribution followed by sampling and sorting $n$ i.i.d. points $t^{k}\sim q$. To draw a sample from our coupling $\rho$, we first sample a data sequence $\gamma_{1}\sim\mu_{1}$, followed by sampling $N(\gamma_{1})$ events independently from $q$ and sorting to produce a draw $\gamma_{0}\sim\mu_{0}$.
Interpolant Construction
We now proceed to construct our interpolant $\eta_{s}\in\mathbb{P}(\Gamma)$. We will construct this path of TPPs via a local procedure which we then marginalize over a given balanced coupling. Here, we extend flow matching techniques (Lipman et al., 2023; Tong et al., 2024) to design our sequence-level interpolants, but we emphasize that this is only possible under a balanced coupling as the number of events is fixed. To that end, let $\rho$ be any balanced coupling of the reference measure $\mu_{0}$ and the data measure $\mu_{1}$, and suppose $z:=(\gamma_{0},\gamma_{1})\sim\rho$ is a draw from this coupling. As $\rho$ is balanced, we have $\gamma_{0}=\sum_{k=1}^{n}\delta[t_{0}^{k}]$ and $\gamma_{1}=\sum_{k=1}^{n}\delta[t_{1}^{k}]$ are both a collection of $n$ events. As TPPs are fully characterized by their joint distributions over event times, we will henceforth describe our procedure for a fixed (but arbitrary) number of events $n$. First, we the define measure $\gamma_{s}^{z}\in\Gamma$ via
@@
We now lift this deterministic path $(\gamma_{s}^{z})_{s=0}^{1}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s=0}^{1}\subset\mathbb{P}(\Gamma)$. We define the point process distribution $\eta_{s}^{z}\in\mathbb{P}(\Gamma)$ implicitly by adding independent Gaussian noise to each of the events in $\gamma_{s}^{z}$. That is, a draw $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ may be simulated via
@@
Finally, for $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}\in\mathbb{P}(\Gamma)$ via $\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$. Observe that, by construction, the event count distribution $\eta_{s}(n)$ is given by $\mu_{1}(n)$ for all $s\in[0,1]$. This path of TPP distributions $\eta_{s}$ approximately interpolates from the reference TPP $\mu_{0}$ at $s=0$ to the data TPP $\mu_{1}$ at $s=1$, in the sense that at the endpoints, the joint event time distributions $\eta_{0}^{n}(t^{1},\dots,t^{n})$ and $\eta_{1}^{n}(t^{1},\dots,t^{n})$ are given by a convolution of $\mu_{0}^{n}(t^{1},\dots,t^{n})$ and $\mu_{1}^{n}(t^{1},\dots,t^{n})$ with the Gaussian $\mathcal{N}(0,\sigma^{2}I_{n})$. As $\sigma^{2}\downarrow 0$, it is clear that we recover a genuine interpolant (Tong et al., 2024).
For single event $t_{s}^{k}$, we have constructed a path of distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$. This transport of a Gaussian can be achieved infinitesimally through the vector field $v_{s}^{k}(t)=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024). Thus, the evolution (4) is generated by the vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$,